# Evacuation Planning on Time-Expanded Networks with Integrated Wildfire Information

### v1

@@ 4.2 Construction of an Initial Evacuation Plan
If some sink is reachable, we proceed to construct the WTEN; see Section 3. The addition of a super source and super sink sets up a classical maximum flow problem. For the solution of (WTEN LP), we use NetworkX’s built-in maximum flow algorithm. It allows the user to choose which algorithm they wish to use, such as Edmonds and Karps [13] or Goldberg and Tarjan’s [17]. For a complete list of options we refer the reader to [20]. We use Dinic’s shortest augmenting path algorithm [11] to solve (WTEN LP). As we will see in Section 5.4, the solution of the problem is a small and efficient part of the overall running time of our algorithm.
@@
Let (LABEL:Dynamic) represent a dynamic flow problem on a dynamic network $G=(N,A,T)$ and let $F(t)$ be a fire set for $t\in[T]_{0}$. The associated (WTEN LP) can be set up and solved in time $O(T|A|\cdot(T|N|+f_{A}))$ in the arithmetic model of computation.
@@
Further, $G_{W}$ satisfies $|N_{W}|\leq(T+1)|N|+2\in O(T|N|)$ and $|A_{W}|\leq T(|A|+|N|)+|N|\in O(T|A|)$, by Lemma 1. A running time of $O(|A_{W}|\cdot|N_{W}|)\subset O(T|A|\cdot T|N|)$ for the solution of a maximum flow problem is possible using the algorithm in [46]. Together, this gives the claimed running time.
We note that the software package used contains simpler maximum flow algorithms than the one in [46]. The options mentioned above have running times of $O(T|N|\cdot(T|A|)^{2})$ for Edmonds-Karp, $O(T|N|^{3})$ for Goldberg and Tarjan’s algorithm, and $O((T|N|)^{2}\cdot T|A|)$ for Dinic’s shortest augmenting path algorithm.
@@
For a fixed time horizon $T$, by construction, a solution of (WTEN LP) is equivalent to solving the associated dynamic flow problem. In the WTEN, the cut $(\hat{s},N_{W}\backslash\{\hat{s}\})$ (between super source and the rest of the network) has capacity $u_{\max}=\sum_{j\in S_{W}}u_{\hat{s}j}$ equivalent to the number of people to evacuate (a complete evacuation), and so a maximum flow value for the WTEN is bounded above by $u_{\max}$.
@@
Let $T_{0}=\sum_{ij\in A}\lambda_{ij}$ be the sum of travel times for all arcs in the original dynamic network. Observe that the travel time for any path from any node to another in the original network is bounded above by $T_{0}$; in particular, this includes paths from a source to a sink. Thus, if any sink can be reached from some source, (WTEN LP) has a feasible solution for WTEN $G_{W}=(N_{W},A_{W})$ with time horizon $T_{0}$ that contains flow along a path in the original network.
@@
Input new fire data and $T$ from initial plan Construct new fire for $t_{\text{fire}}\leq t<T$ Combine with old fire from $t_{\text{reopt}}\leq t<t_{\text{fire}}$ Everyone evacuated? $T=T+1$ $T>T_{\text{max}}$ $T=T_{\text{sol}}$ Return maximum flow solution for time horizon $T$ Solve (WTEN LP) Construct WTEN for $t_{\text{reopt}}\leq t\leq T$ noyesyesno
@@ 4.3 Real-time Solution Updates
In the above, we described the setup of WTEN and the computation of an initial feasible solution to (WTEN LP) at minimal time horizon $T$ via a classical maximum flow algorithm. This feasible solution is the initial evacuation plan for an area.
@@
The shortest augmenting path algorithm is then used to solve (WTEN LP) for this smaller, new WTEN to obtain an updated evacuation plan. Figure 7(a) and 7(b) show an example initial and updated evacuation plan. According to new fire information, nodes 7 and 11 are going to be overtaken and thus removed from the network. Arcs (1,4), (4,8), (4,9), (8,15), (9,12) and (12,15) are now dangerously close to the fire and have reduced capacity (and in turn flow). In this example, a maximum flow value towards sink node 15 decreases; the change in fire may lead to a scenario in which the evacuation is not completed within the same time horizon.
@@
The (WTEN LP) with fixed flows $x_{ij}(t)$ for $t<t_{\text{reopt}}$ and based on $F^{\prime}(t)$ can be set up and solved in time $O(T|A|\cdot(T|N|+f_{A}))$ in the arithmetic model of computation.
@@
$S^{\prime}_{W}=\{j(t^{\prime})\in N_{W}:(i(t),j(t^{\prime}))\in A_{W},x_{i(t)j(t^{\prime})}>0,t^{\prime}=t+\lambda_{i(t)j(t^{\prime})},t<t_{\text{reopt}}\}$
@@
$\sum\limits_{t=0}^{t_{\text{reopt}}-1}\sum\limits_{(i(t),j(t^{\prime}))\in A_{W}}x_{i(t)j(t^{\prime})}.$
One then replaces the number of time instances $|[T+1]_{0}|$ with $|[T+1-t_{\text{reopt}}]_{0}|$, shifting all information for time instances $t\geq t_{\text{reopt}}$ to time instances $t-t_{\text{reopt}}$ instead, and solve the new, smaller (WTEN LP).
@@ 5 Illustrative Applications
For each case, we place an approximated fire in either a historic fire location or, in the case of Lyons, Colorado, in an area of high fire danger [47]. We begin by constructing an initial evacuation plan based on these fire locations. Then, we create a change (or multiple changes) in the fire at a specified time instance. Finally, we update the initial evacuation plan to account for the change. For case 2 (Butte County), we focus on an evacuation of the town of Paradise, California. An initial evacuation plan is run over the large Butte County network, while the update portion of our approach is run only on a network for the town itself. This lets us discuss scalability and computation times for the two vastly different problem sizes.
@@ 5.1 Case 1: Lyons, Colorado
The first location is Lyons, Colorado, a town located at the base of the foothills of Colorado with few roads in and out of the town. The fire danger in this town is high due to the location [47]. We chose to place the fire near the northeastern part of the town, which is a large area of open land and is under considerable fire watch. The sources and sinks, including their supply/demand values, are based on the population of the town and showcase a difference between the initial and updated evacuation plans.
We give each of the two source nodes a supply of 4,000 people. Lyons has a population of roughly 6,000 people [47]. This overestimation in supply is done to visualize an aversion to fire risk in the evacuation routes chosen. Everyone is able to evacuate with a total flow value of 8,000 within 24 minutes.
@@ 5.2 Case 2: Butte County, California
Butte County California is the location of the deadliest wildfire in California’s history: the Camp Fire [39]. While the Camp Fire mainly affected the residents of Paradise, California, we elect to use the entire road network for Butte County for an initial evacuation plan to provide an example on a large network; see Table 1. The road network used has a coarser tolerance of 50 meters in order to keep scale and running times reasonable, while still maintaining enough detail to accurately capture the evacuation. The fire location is based on the Camp Fire. The sources are nodes in Paradise, California, while the sinks lie outside of the town. The scenario is set up to resemble the situation during the Camp Fire evacuation.
@@ 5.3 Case 3: Alexandroupolis, Greece
In 2023, a series of fires in Greece broke out in a 24-hour time frame, burning a total of roughly 730 square kilometers. It is one of the largest wildfires in European history [52]. Alexandroupolis, Greece was one of the cities affected. Our estimated fire locations are based on this historic fire, with sources selected as nodes close to the fire, in most imminent need of evacuation. A single sink is chosen in the safe south of the city. The supply values are divided among three sources as 1,000, 1,500, and 2,500. A complete evacuation is possible in 48 minutes.
@@ 6 Final Remarks
There are several natural directions for future work. The current black box methods used for distance computations between roads and fire sets are a bottleneck in overall algorithm running time. It would be beneficial to explore creating in-house methods for these computations, either exact or approximate. One could also test the impact of data extrapolation, i.e., reusing distance information over a certain number of time instances before recomputing. Other algorithmic improvements to consider include warm starting a combinatorial algorithm for the reoptimization of (WTEN LP) (using new fire data). This would be based on the initial dual optimal solution and could further decrease the time needed to construct an updated plan.
Further challenges in evacuating planning include increasing evacuation compliance and ensuring equity [58]. Many measures of equity have been considered within a mathematical optimization model, however, a major hurdle to overcome in all cases is the resulting increase in computational time. An important question that merits further research is how to best account for equity in evacuation planning without sacrificing model efficiency. A starting point for this could be to incorporate equity measures such as the Kolm Pollak Equally Distributed Equivalent which has been shown to scale well to large problem instances in related settings [25, 28, 38]. One could also consider first introducing equity measures into one key step of the process, for example, the initial assignment portion of the algorithm described in [32].
@@ Acknowledgments
This work was partially supported by Air Force Office of Scientific Research grant FA9550-21-1-0233 (Complex Networks) and NSF grant 2006183 (Algorithmic Foundations, Division of Computing and Communication Foundations).

### v2

@@ 4.2 Construction of an Initial Evacuation Plan
If some sink is reachable, we proceed to construct the WTEN; see Section 3. The addition of a super source and super sink sets up a classical maximum flow problem. For the solution of (LABEL:WTEN_LP), we use NetworkX’s built-in maximum flow algorithm. It allows the user to choose which algorithm they wish to use, such as Edmonds and Karp’s maximum flow algorithm (Edmonds and Karp, 1972) or Goldberg and Tarjan’s push relabel algorithm (Goldberg and Tarjan, 1988). For a complete list of options we refer the reader to (Hagberg et al., 2008). We use Dinic’s shortest augmenting path algorithm (Dinic, 1970) to solve (LABEL:WTEN_LP). As we will see in Section 5.4, the solution of the problem is a small and efficient part of the overall running time of our algorithm.
@@
Let (LABEL:Dynamic) represent a dynamic flow problem on a dynamic network $G=(N,A,T)$ and let $F(t)$ be a fire set for $t\in[T]_{0}$. The associated (LABEL:WTEN_LP) can be set up and solved in time $O(T|A|\cdot(T|N|+f_{A}))$ in the arithmetic model of computation.
@@
Further, $G_{W}$ satisfies $|N_{W}|\leq(T+1)|N|+2\in O(T|N|)$ and $|A_{W}|\leq T(|A|+|N|)+|N|\in O(T|A|)$, by Lemma 1. A running time of $O(|A_{W}|\cdot|N_{W}|)\subset O(T|A|\cdot T|N|)$ for the solution of a maximum flow problem is possible using the algorithm in (Orlin, 2013). Together, this gives the claimed running time.
We note that the software package used contains simpler maximum flow algorithms than the one in (Orlin, 2013). The options mentioned above have running times of $O(T|N|\cdot(T|A|)^{2})$ for Edmonds-Karp, $O(T|N|^{3})$ for Goldberg and Tarjan’s algorithm, and $O((T|N|)^{2}\cdot T|A|)$ for Dinic’s shortest augmenting path algorithm.
@@
For a fixed time horizon $T$, by construction, a solution of (LABEL:WTEN_LP) is equivalent to solving the associated dynamic flow problem. In the WTEN, the cut $(\hat{s},N_{W}\backslash\{\hat{s}\})$ (between super source and the rest of the network) has capacity $u_{\max}=\sum_{j\in S_{W}}u_{\hat{s}j}$ equivalent to the number of people to evacuate (a complete evacuation), and so a maximum flow value for the WTEN is bounded above by $u_{\max}$.
@@
Let $T_{0}=\sum_{ij\in A}\lambda_{ij}$ be the sum of travel times for all arcs in the original dynamic network. Observe that the travel time for any path from any node to another in the original network is bounded above by $T_{0}$; in particular, this includes paths from a source to a sink. Thus, if any sink can be reached from some source, (LABEL:WTEN_LP) has a feasible solution for WTEN $G_{W}=(N_{W},A_{W})$ with time horizon $T_{0}$ that contains flow along a path in the original network.
@@
In this subsection, we presented the procedure our algorithm uses to create an initial evacuation plan before a fire occurs. This could be combined with the use of one or more predictive fire models, enabling an expert to determine initial evacuation routes under a variety of events. In this way, our tool can be used to test and develop many different evacuation scenarios that can be shared with residents and practitioners ahead of time. The next part of the algorithm, and the next subsection, focuses on real-time or short notice updates.
Input new fire data and $T$ from initial plan Construct new fire for $t_{\text{fire}}\leq t<T$ Combine with old fire from $t_{\text{reopt}}\leq t<t_{\text{fire}}$ Everyone evacuated? $T=T+1$ $T>T_{\text{max}}$ $T=T_{\text{sol}}$ Return maximum flow solution for time horizon $T$ Solve (LABEL:WTEN_LP) Construct WTEN for $t_{\text{reopt}}\leq t\leq T$ noyesyesno
@@ 4.3 Real-time Solution Updates
In the above, we described the setup of WTEN and the computation of an initial feasible solution to (LABEL:WTEN_LP) at minimal time horizon $T$ via a classical maximum flow algorithm. This feasible solution is the initial evacuation plan for an area.
@@
The shortest augmenting path algorithm is then used to solve (LABEL:WTEN_LP) for this smaller, new WTEN to obtain an updated evacuation plan. Figure 7(a) and 7(b) show an example initial and updated evacuation plan. According to new fire information, nodes 7 and 11 are going to be overtaken and thus removed from the network. Arcs (1,4), (4,8), (4,9), (8,15), (9,12) and (12,15) are now dangerously close to the fire and have reduced capacity (and in turn flow). In this example, a maximum flow value towards sink node 15 decreases; the change in fire may lead to a scenario in which the evacuation is not completed within the same time horizon.
@@
The (LABEL:WTEN_LP) with fixed flows $x_{ij}(t)$ for $t<t_{\text{reopt}}$ and based on $F^{\prime}(t)$ can be set up and solved in time $O(T|A|\cdot(T|N|+f_{A}))$ in the arithmetic model of computation.
@@
$S^{\prime}_{W}=\{j(t^{\prime})\in N_{W}:(i(t),j(t^{\prime}))\in A_{W},x_{i(t)j(t^{\prime})}>0,t^{\prime}=t+\lambda_{i(t)j(t^{\prime})},t<t_{\text{reopt}}\}$ (36)
@@
$\sum\limits_{t=0}^{t_{\text{reopt}}-1}\sum\limits_{(i(t),j(t^{\prime}))\in A_{W}}x_{i(t)j(t^{\prime})}.$ (37)
One then replaces the number of time instances $|[T+1]_{0}|$ with $|[T+1-t_{\text{reopt}}]_{0}|$, shifting all information for time instances $t\geq t_{\text{reopt}}$ to time instances $t-t_{\text{reopt}}$ instead, and solve the new, smaller (LABEL:WTEN_LP).
@@ 5 Illustrative Applications
For each case, we place an approximated fire in either a historic fire location or, in the case of Lyons, Colorado, in an area of high fire danger (Owen and Hoffman, 2011). We begin by constructing an initial evacuation plan based on these fire locations. Then, we create a change (or multiple changes) in the fire at a specified time instance. Finally, we update the initial evacuation plan to account for the change. For case 2 (Butte County), we focus on an evacuation of the town of Paradise, California. An initial evacuation plan is run over the large Butte County network, while the update portion of our approach is run only on a network for the town itself. This lets us discuss scalability and computation times for the two vastly different problem sizes.
@@ 5.1 Case 1: Lyons, Colorado
The first location is Lyons, Colorado, a town located at the base of the foothills of Colorado with few roads in and out of the town. The fire danger in this town is high due to the location (Owen and Hoffman, 2011). We chose to place the fire near the northeastern part of the town, which is a large area of open land and is under considerable fire watch. The sources and sinks, including their supply/demand values, are based on the population of the town and showcase a difference between the initial and updated evacuation plans.
We give each of the two source nodes a supply of 4,000 people. Lyons has a population of roughly 6,000 people (Owen and Hoffman, 2011). This overestimation in supply is done to visualize an aversion to fire risk in the evacuation routes chosen. Everyone is able to evacuate with a total flow value of 8,000 within 24 minutes.
@@ 5.2 Case 2: Butte County, California
Butte County California is the location of the deadliest wildfire in California’s history: the Camp Fire (Maranghides et al., 2023). While the Camp Fire mainly affected the residents of Paradise, California, we elect to use the entire road network for Butte County for an initial evacuation plan to provide an example on a large network; see Table 1. The road network used has a coarser tolerance of 50 meters in order to keep scale and running times reasonable, while still maintaining enough detail to accurately capture the evacuation. The fire location is based on the Camp Fire. The sources are nodes in Paradise, California, while the sinks lie outside of the town. The scenario is set up to resemble the situation during the Camp Fire evacuation.
@@ 5.3 Case 3: Alexandroupolis, Greece
In 2023, a series of fires in Greece broke out in a 24-hour time frame, burning a total of roughly 730 square kilometers. It is one of the largest wildfires in European history (San-Miguel-Ayanz et al., 2024). Alexandroupolis, Greece was one of the cities affected. Our estimated fire locations are based on this historic fire, with sources selected as nodes close to the fire, in most imminent need of evacuation. A single sink is chosen in the safe south of the city. The supply values are divided among three sources as 1,000, 1,500, and 2,500. A complete evacuation is possible in 48 minutes.
@@ 6 Final Remarks
There are several natural directions for future work. The current black box methods used for distance computations between roads and fire sets are a bottleneck in overall algorithm running time. It would be beneficial to explore creating in-house methods for these computations, either exact or approximate. One could also test the impact of data extrapolation, i.e., reusing distance information over a certain number of time instances before recomputing. Other algorithmic improvements to consider include warm starting a combinatorial algorithm for the reoptimization of (LABEL:WTEN_LP) (using new fire data). This would be based on the initial dual optimal solution and could further decrease the time needed to construct an updated plan.
Further challenges in evacuating planning include increasing evacuation compliance and ensuring equity (Zhao and Wong, 2021). Many measures of equity have been considered within a mathematical optimization model, however, a major hurdle to overcome in all cases is the resulting increase in computational time. An important question that merits further research is how to best account for equity in evacuation planning without sacrificing model efficiency. A starting point for this could be to incorporate equity measures such as the Kolm Pollak Equally Distributed Equivalent which has been shown to scale well to large problem instances in related settings (Horton et al., 2024; Kolm, 1976; Maguire and Sheriff, 2020). One could also consider first introducing equity measures into one key step of the process, for example, the initial assignment portion of the algorithm described in (Lim et al., 2012).
@@ Acknowledgments
This work was supported in part by the Air Force Office of Scientific Research [grants FA9550-21-1-0233 and FA9550-24-1-0240] (Complex Networks); and the National Science Foundation [grant 2006183] (Algorithmic Foundations, Division of Computing and Communication Foundations).